Write 5(1.06^2x+1)=11 as a logarithmic equation then solve
Answer by lwsshak3(11628)

Write 5(1.06^2x+1)=11 as a logarithmic equation then solve
**
5(1.06^2x+1)=11
=5*1.06^2x+5=11
=5*1.06^2x=11-5=6
=1.06^2x=6/5
Take log of both sides
2xlog(1.06)=log(6/5)
2x=log(6/5)/log(1.06)=3.129
x=1.5645
Check:
5(1.06^2x+1)=5*(1.06^(2*1.5645)+1)=5*2.2=11